# Linear Complexity Gibbs Sampling for Generalized Labeled Multi-Bernoulli Filtering

### v1

@@ Abstract
Generalized Labeled Multi-Bernoulli (GLMB) densities arise in a host of multi-object system applications analogous to Gaussians in single-object filtering. However, computing the GLMB filtering density requires solving NP-hard problems. To alleviate this computational bottleneck, we develop a linear complexity Gibbs sampling framework for GLMB density computation. Specifically, we propose a tempered Gibbs sampler that exploits the structure of the GLMB filtering density to achieve an $\mathcal{O}(T(P+M))$ complexity, where $T$ is the number of iterations of the algorithm, $P$ and $M$ are the number hypothesized objects and measurements. This innovation enables an $\mathcal{O}(T(P+M+\log(T))+PM)$ complexity implementation of the GLMB filter. Convergence of the proposed Gibbs sampler is established and numerical studies are presented to validate the proposed GLMB filter implementation.
@@
Random finite sets, Multi-object tracking, Generalized labeled multi-Bernoulli, Tempered Gibbs sampling
@@ I Introduction
The aim of multi-object tracking (MOT) is to estimate the number of objects and their trajectories from noisy sensor data. The challenges are the unknown and time-varying number of objects, accompanied by false alarms, misdetections, and data association uncertainty, culminating in computational bottlenecks for most real world applications. Notwithstanding this, numerous solutions have been developed, with multiple hypothesis tracking [1], joint probabilistic data association [2], and random finite set (RFS) [3] being the most widely used approaches. The RFS approach, in particular, is gaining substantial interest due to its versatile multi-object state space models [4], and efficient solutions [5]. Various tractable RFS multi-object filters have been devised ranging from the probability hypothesis density localization filters [4], [6] to the generalized labeled multi-Bernoulli (GLMB) tracking filter [7].
@@
The GLMB filter [7] is an analytic solution to the RFS multi-object filter that provides tracks and their provisional identities/labels to meet MOT requirements [1]. The unique feature of this formulation is the provision for principled trajectory estimation (even with only single-scan filtering), capable of tracking millions of objects, online [5]. Moreover, the GLMB family of densities is furnished with elegant mathematical properties, offering a versatile set of tools including, conjugacy [7], closure under truncation with analytic truncation error [8], analytic approximation of general labeled multi-object density with minimal Kullback-Leibler divergence [9], analytic Cauchy-Schwarz divergence and void probabilities [10]. These properties enabled MOT with multiple sensors and scans [8, 11, 12], non-standard models [9, 13, 14, 15], unknown system parameters [16], and distributed implementations [17, 18, 19, 20], as well as multi-object control solutions [10, 21]. Moreover, the GLMB filter has found a host of applications from robotics [22], sensor networks [23, 24], cell biology [14, 25] to audio/video processing [15, 26].
@@
The main computational bottleneck in GLMB filtering is the truncation of the multi-object filtering density [7]. Truncation by discarding terms with small weights minimizes the $L_{1}$-error [8], and can be posed as a ranked assignment problem, solvable by Murty’s algorithm [27] with cubic complexity in both the number of measurements and hypothesized objects [28, 29]. A more efficient solution based on Gibbs sampling (GS) was proposed in [30], which incurs an $\mathcal{O}(TP^{2}M)$ complexity, where $T$, the number of iterates of the algorithm, dominates the number of hypothesized objects and measurements $P$ and $M$. GS is an efficient Markov Chain Monte Carlo technique for sampling from complex probability distributions, popularized by the seminal work of Geman and Geman [31], which opened up applications in many disciplines ranging from statistics, engineering to computer science, and is still an active research topic.
@@
Following the strategy of selecting significant GLMB components by random sampling [30], a number of GLMB truncation techniques have been developed. In [32], an approximate GLMB filter with linear complexity in the number of hypothesized objects was proposed by neglecting the standard data association requirement of at most one measurement per object. While this technique can be modified to accommodate the data association requirement, the complexity reverts to quadratic in number of hypothesized objects [32]. In [33], a herded Gibbs sampling implementation of the labeled multi-Bernoulli (LMB) filter [34]–a one-term approximation of the GLMB filter–was developed. However, this implementation is slower than the systematic-scan Gibbs sampling (SGS) GLMB implementation [30], not to mention that the (exact) LMB filter suffers from track fragmentation and track switching. In [35] and [36], the cross-entropy method [37] was applied, respectively, to multi-sensor GLMB filtering and its distributed version, but with higher complexity than the SGS implementation in [11]. One effective technique to reduce computation times is the use of spatial search to decompose the filtering density into independent GLMBs that can be computed in parallel [5]. Nonetheless, the time complexity of the independent GLMB filters has not been reduced.
@@
To alleviate the computational bottleneck in GLMB filtering, this paper proposes a tempered Gibbs sampling (TGS) framework for selecting significant GLMB components with linear complexity, i.e., $\mathcal{O}(T(P+M))$. Similar to the widely known random-scan Gibbs sampling (RGS), TGS randomly selects a coordinate to update, but provides the mechanism to improve mixing and sample diversity [38]. However, generic TGS incurs an $\mathcal{O}(TP^{2}M)$ complexity. To this end we develop an innovative decomposition of the conditionals that enables TGS to be reduced to $\mathcal{O}(T(P+M))$ complexity, with negligible additional memory (compared to generic $\mathcal{O}(TP^{2}M)$ TGS). Convergence and statistical efficiency of the proposed TGS algorithm are also discussed. Note that our solution is not only restricted to GLMB truncation, but amenable to a wide range of applications of the ranked assignment problem [39].
@@
The proposed TGS framework enables RGS to be implemented with $\mathcal{O}(T(P+M))$ complexity, whereas generic RGS incurs $\mathcal{O}(TPM)$. Further, we propose deterministic-scan Gibbs sampling (DGS), an efficient algorithm using deterministic coordinate selection, as opposed to RGS’s completely random coordinate selection. DGS’s better mixing and sample diversity compared to RGS is validated by numerical studies. In addition, using DGS, we show how SGS can be implemented with $\mathcal{O}(TPM)$ complexity. We also present numerical studies in MOT to discuss the trade-offs between tracking performance and computational load in the proposed TGS framework. Due to the computation of the cost matrix, the resultant GLMB filter implementation incurs an additional $\mathcal{O}(PM)$ complexity.
@@ II Preliminaries
$x_{m:n}$ $x_{m},x_{m\text{+}1},\ldots,x_{n}$
@@
$1_{S}(\cdot)$ Indicator function for a given set $S$
$\delta_{X}[Y]$ Kronecker-$\delta$, $\delta_{X}[Y]=1$ if $X=Y$, 0 otherwise
$\mathcal{F}(S)$ Class of all finite subsets of a given set $S$
[…]
$|X|$ Cardinality of a set $X$
@@
$\mathcal{L}(\boldsymbol{x})$ Label of a (labeled) state $\boldsymbol{x}$, $\mathcal{L}((x,\ell))=\ell$
$\mathcal{L}(\boldsymbol{X})$ Labels of a (labeled) set $\boldsymbol{X}$, $\{\mathcal{L}(\boldsymbol{x}):\boldsymbol{x}\in\boldsymbol{X}\}$
$\Delta(\boldsymbol{X})$ Distinct label indicator $\delta_{|\boldsymbol{X}|}(|\mathcal{L}(\boldsymbol{X})|)$
$\gamma$ Positive 1-1 association map
@@
$\alpha$ Mixture parameter of proposal
$\beta$ Tempering parameter of proposal
This section presents the necessary background for the development of the main result of this article. We first outline the basics of generalized labeled multi-Bernoulli (GLMB) filtering in Subsection II-A, and then summarize the Gibbs sampling (GS) approach to GLMB truncation in Subsection II-B. Frequently used notations are summarized in Table I.
@@ II-A1 Multi-object State
In GLMB filtering, the system state to be estimated at each time is the set $\boldsymbol{X}$ of labeled states of the underlying objects, called the multi-object state [7]. Each labeled single-object state $\boldsymbol{x}\in\boldsymbol{X}$ is an ordered pair $(x,\ell)$ in the product space $\mathbb{\mathbb{X}\times\mathbb{L}}$, where $\mathbb{X}$ is a (finite dimensional) state space, and $\mathbb{L}$ is a discrete space called the label space. Let $\mathbb{B}_{\tau}$ denote the label space of objects born at time $\tau$, then the space $\mathbb{L}$ of all labels up to time $k$ is given by the disjoint union $\uplus_{\tau\texttt{=}0}^{k}\mathbb{B}_{\tau}$. The state $x$ of an object varies with time, while its label or identity $\ell$ (usually consists of the object’s time of birth and an index to distinguish those born at the same time) is time-invariant. The trajectory of an object is a sequence of consecutive labeled states with a common label.
The cardinality $\left|\boldsymbol{X}\right|$ (i.e., number of elements) of the multi-object state $\boldsymbol{X}$ varies with time due to the appearance and disappearance of objects. In addition, a multi-object state $\boldsymbol{X}$ at any time must have distinct labels. More concisely, let $\mathcal{L}\left(\boldsymbol{x}\right)$ denote the label of $\boldsymbol{x}$, and for any finite $\boldsymbol{X}\subset\mathbb{X}\times\mathbb{L}$, define the label set $\mathcal{L}\left(\boldsymbol{X}\right)\triangleq\left\{\mathcal{L}\left(\boldsymbol{x}\right):\boldsymbol{x}\in\boldsymbol{X}\right\}$ and distinct label indicator $\Delta(\boldsymbol{X})\triangleq\delta_{|\boldsymbol{X}|}\big{[}|\mathcal{L}(\boldsymbol{X})|\big{]}$. Then for $\boldsymbol{X}$ to be valid multi-object state, we require $\Delta\left(\boldsymbol{X}\right)=1$.
In Bayesian estimation, the single-object state and the measurement are modeled as random variables. Hence, for multi-object estimation the multi-object state and measurement are modeled as Random Finite Sets (RFSs). In the following we describe the so-called standard multi-object models for the dynamics and observations in a multi-object system.
@@ II-A3 Multi-object Observation
Each element $\boldsymbol{x}\in\boldsymbol{X}$ is either detected with probability $P_{D}(\boldsymbol{x})$ and generates an observation $z$ at the sensor with likelihood $g(z|\boldsymbol{x})$, or misdetected with probability $1-P_{D}(\boldsymbol{x})$. In addition to the detections, the sensor also receives clutter, modeled by a Poisson RFS with intensity function $\kappa$. The multi-object observation $Z$ is the union of detections and clutter. It is assumed that conditional on the multi-object state, objects are detected independently from each other and that clutter and detections are independent.
The association of objects with sensor measurements (at time $k$) is described by a positive 1-1 mapping $\gamma\!:\!\mathbb{L}\!\rightarrow\!\{\texttt{\text{-}}1{\textstyle:}|Z|\}$, i.e., a mapping where no two distinct arguments are mapped to the same positive value [7]. The association map $\gamma$ specifies that object (with label) $\ell$ generates measurement $z_{\gamma(\ell)}\in Z$, with $\gamma(\ell)=0$ if it is undetected, and $\gamma(\ell)=-1$ if it does not exist. The positive 1-1 property ensures each measurement comes from at most one object. Let $\mathcal{L}(\gamma)\triangleq\{\ell\in\mathbb{L}:\gamma(\ell)\geq 0\}$ denote the set of live labels [^1] Note the distinction from $\mathcal{L}(\boldsymbol{X})$, the labels of a labeled set. of $\gamma$, and $\Gamma$ denote the space of all association maps, then the multi-object likelihood function can be written as [11]
$g(Z|$ $\boldsymbol{X})\propto\underset{\gamma\in\Gamma}{\sum}\delta_{\mathcal{L}(\gamma)}[\mathcal{L}(\boldsymbol{X})][\psi_{Z}^{(\gamma\circ\mathcal{L}(\cdot))}(\cdot)]^{\boldsymbol{X}},$ (1)
@@
$\psi_{\{z_{1:M}\}}^{(j)}(\boldsymbol{x})=\begin{cases}\frac{P_{D}(\boldsymbol{x})g(z_{j}|\boldsymbol{x})}{\kappa(z_{j})},&j>0\\ 1-P_{D}(\boldsymbol{x}),&j=0\end{cases}.$ (2)

### v2

@@ Abstract
Generalized Labeled Multi-Bernoulli (GLMB) densities arise in a host of multi-object system applications analogous to Gaussians in single-object filtering. However, computing the GLMB filtering density requires solving NP-hard problems. To alleviate this computational bottleneck, we develop a linear complexity Gibbs sampling framework for GLMB density computation. Specifically, we propose a tempered Gibbs sampler that exploits the structure of the GLMB filtering density to achieve an $\mathcal{O}(T(P+M))$ complexity, where $T$ is the number of iterations of the algorithm, $P$ and $M$ are the number hypothesized objects and measurements. This innovation enables the GLMB filter implementation to be reduced from an $\mathcal{O}(TP^{2}M)$ complexity to $\mathcal{O}(T(P+M+\log T)+PM)$. Moreover, the proposed framework provides the flexibility for trade-offs between tracking performance and computational load. Convergence of the proposed Gibbs sampler is established, and numerical studies are presented to validate the proposed GLMB filter implementation.
@@
Random finite sets, Multi-object tracking, Generalized labeled multi-Bernoulli, Tempered Gibbs sampling.
@@ I Introduction
The aim of multi-object tracking (MOT) is to estimate the number of objects and their trajectories from noisy sensor data. The challenges are the unknown and time-varying number of objects, accompanied by false alarms, misdetections, and data association uncertainty, culminating in computational bottlenecks for most real world applications. Notwithstanding this, numerous solutions have been developed, with multiple hypothesis tracking [1], joint probabilistic data association [2], and random finite set (RFS) [3] being the most widely used approaches. The RFS approach, in particular, is gaining substantial interest due to its versatile multi-object state space models [4], and efficient solutions [5]. Various tractable RFS multi-object filters have been devised ranging from the probability hypothesis density localization filters [4, 6] to the generalized labeled multi-Bernoulli (GLMB) tracking filter [7].
@@
The GLMB filter [7] is an analytic solution to the RFS multi-object filter that provides tracks and their provisional identities/labels to meet MOT requirements [1]. The unique feature of this formulation is the provision for principled trajectory estimation (even with only single-scan filtering), capable of tracking millions of objects, online [5]. Moreover, the GLMB family of densities is furnished with elegant mathematical properties, offering a versatile set of tools including, conjugacy [7], closure under truncation with analytic truncation error [8], analytic approximation of general labeled multi-object density with minimal Kullback-Leibler divergence [9], analytic Cauchy-Schwarz divergence and void probabilities [10]. These properties enabled MOT with multiple sensors and scans [8, 11, 12], non-standard models [9, 13, 14, 15], unknown system parameters [16], multi-object control solutions [10, 17], as well as distributed implementations [18, 19, 20]. The GLMB filter is also amenable to parallelization that reduces computation times. For instance, in [5], spatial search was used to decompose the filtering density into independent GLMBs that are processed in parallel, while, in [21], a parallel centralized implementation for multiple sensors was developed. The GLMB filter has also found a host of applications from robotics [22], sensor networks [23, 24], cell biology [14, 25] to audio/video processing [15, 26].
@@
The main computational bottleneck in GLMB filtering is the truncation of the multi-object filtering density [7]. Truncation by discarding terms with small weights minimizes the $L_{1}$-error [8], and can be posed as a ranked assignment problem, solvable by Murty’s algorithm [27] with cubic complexity in both the number of measurements and hypothesized objects [28, 29]. A more efficient solution based on Gibbs sampling (GS) was proposed in [30], which incurs an $\mathcal{O}(TP^{2}M)$ complexity, where $T$, the number of iterates of the algorithm, dominates the number of hypothesized objects and measurements $P$ and $M$. GS is an efficient Markov Chain Monte Carlo (MCMC) technique for sampling from complex probability distributions, popularized by the seminal work of Geman and Geman [31], which opened up applications in many disciplines ranging from statistics, engineering to computer science, and is still an active research topic.
@@
Following the strategy of selecting significant GLMB components by random sampling [30], a number of GLMB truncation techniques have been developed. In [32], an approximate GLMB filter with linear complexity in the number of hypothesized objects was proposed by neglecting the standard data association requirement of at most one measurement per object. While this technique can be modified to accommodate the data association requirement, the complexity reverts to quadratic in number of hypothesized objects [32]. In [33], a herded Gibbs sampling implementation of the labeled multi-Bernoulli (LMB) filter [34]–a one-term approximation of the GLMB filter–was developed. However, this implementation is slower than the systematic-scan Gibbs sampling (SGS) GLMB implementation [30], not to mention that, compared to the (exact) GLMB filter, the LMB filter is more prone to track fragmentation and track switching. In [35] and [36], the cross-entropy method [37] was applied, respectively, to multi-sensor GLMB filtering and its distributed version, but with higher complexity than the SGS implementation in [11].
@@
To alleviate the computational bottleneck in GLMB filtering, this paper proposes a tempered Gibbs sampling (TGS) framework for selecting significant GLMB components, with linear complexity, i.e., $\mathcal{O}(T(P+M))$. Similar to the widely known random-scan Gibbs sampling (RGS), TGS randomly selects a coordinate to update, but provides the mechanism to improve mixing and sample diversity [38]. However, generic TGS incurs an $\mathcal{O}(TP^{2}M)$ complexity. To this end, we develop an innovative decomposition of the conditionals that enables TGS to be reduced to $\mathcal{O}(T(P+M))$ complexity, with negligible additional memory (compared to generic $\mathcal{O}(TP^{2}M)$ TGS). The samples generated by the proposed TGS algorithm converge to a tempered distribution (not necessarily the same as the original stationary distribution). Furthermore, keeping in mind that TGS is regarded as the combination of importance sampling and MCMC, the importance-weighted samples indeed converge to the original stationary distribution [38]. The drastic reduction in computational complexity facilitates the development of multi-object control solutions, which require fast computation of the multi-object filtering density for online operations [39, 40, 41, 42]. Moreover, our solution is not only restricted to GLMB truncation, but amenable to a wide range of applications of the ranked assignment problem [43].
@@
The proposed TGS framework enables RGS to be implemented with $\mathcal{O}(T(P+M))$ complexity, whereas generic RGS incurs $\mathcal{O}(TPM)$. Further, we propose deterministic-scan Gibbs sampling (DGS), an efficient algorithm using deterministic coordinate selection, as opposed to RGS’s completely random coordinate selection. DGS’s better mixing and sample diversity compared to RGS are validated by numerical studies. In addition, using DGS, we show how an exact implementation of SGS can be accomplished with $\mathcal{O}(TPM)$ complexity. We also present numerical studies in MOT to discuss the trade-offs between tracking performance and computational load in the proposed TGS framework. Due to the computation of the cost matrix, the resultant GLMB filter implementation incurs an additional $\mathcal{O}(PM)$ complexity. In summary, our main contribution is a TGS framework for GLMB truncation that:
  • Incurs a linear complexity of $\mathcal{O}(T(P+M))$, with negligible additional memory, by exploiting the structure of the GLMB filtering density; and
  • Admits as special cases, RSG, DGS, and SGS implementations, with $\mathcal{O}(T(P+M))$, $\mathcal{O}(TM)$, and $\mathcal{O}(TPM)$ complexities, respectively.
We validate the proposed approach via a series of comprehensive numerical experiments, with considerations for computational efficiency, tracking accuracy, and sample diversity.
@@ II Preliminaries
This section presents the necessary background for the development of the main result of this article. We first outline the basics of generalized labeled multi-Bernoulli (GLMB) filtering in Subsection II-A, and then summarize the Gibbs sampling (GS) approach to GLMB truncation in Subsection II-B. Throughout this paper, single-object states and multi-object states are respectively represented by lower case letters (e.g., $x$) and upper case letters (e.g., $X$). Further, boldfaced symbols denote labeled states or distributions (e.g., $\bm{x}$, $\bm{X}$, $\bm{\pi}$). Frequently used notations are summarized in Table I.
@@
$1_{S}(\cdot)$ Indicator function for a given set $S$
$|X|$ Cardinality of a set $X$
[…]
$\alpha$ Mixture parameter of proposal distribution
$\beta$ Tempering parameter of proposal distribution
$\delta_{X}[Y]$ Kronecker-$\delta$, $\delta_{X}[Y]=1$ if $X=Y$, 00 otherwise
$\Delta(\bm{X})$ Distinct label indicator $\delta_{|\bm{X}|}[|\mathcal{L}(\bm{X})|]$
$\mathcal{F}(S)$ Class of all finite subsets of a given set $S$
@@
$f_{B}(x,\ell)$ Birth probability density of $x$ given label $\ell$
$\Gamma$ Space of all (extended) association map $\gamma$
$\gamma$ (Extended) association map
$\gamma\circ\mathcal{L}(\cdot)$ Function composition $\gamma(\mathcal{L}(\cdot))$
$g(z|\bm{x})$ Single-object likelihood
$g(Z|\bm{X})$ Multi-object likelihood
@@
$\mathcal{L}(\bm{x})$ Label of a (labeled) state $\bm{x}$, $\mathcal{L}((x,\ell))=\ell$
$\mathcal{L}(\bm{X})$ Labels of a (labeled) set $\bm{X}$, $\{\mathcal{L}(\bm{x}):\bm{x}\in\bm{X}\}$
$\mathcal{L}(\gamma)$ Live labels of $\gamma$, $\{\ell\in\mathbb{L}:\gamma(\ell)\geq 1\}$
$M$ Number of measurements (at the next time)
$P$ Number of hypothesized objects
$P_{B}(\ell)$ Birth probability of label $\ell$
$P_{D}(\bm{x})$ Detection probability of labeled state $\bm{x}$
$P_{S}(\bm{x})$ Survival probability of labeled state $\bm{x}$
$\bm{\pi}$ Multi-object filtering density
@@
$T$ Number of iterates
$x_{m:n}$ $x_{m},x_{m\text{+}1},...,x_{n}$
$\xi$ History of association maps
$\Xi$ Some finite discrete space
@@ II-A1 Multi-object State
In GLMB filtering, the system state to be estimated at each time is the set $\bm{X}$ of labeled states of the underlying objects, called the multi-object state [7]. Each labeled single-object state $\bm{x}\in\bm{X}$ is an ordered pair $(x,\ell)$ in the product space $\mathbb{\mathbb{X}\times\mathbb{L}}$, where $\mathbb{X}$ is a (finite dimensional) state space, and $\mathbb{L}$ is a discrete space called the label space. Let $\mathbb{B}_{\tau}$ denote the label space of objects born at time $\tau$, then the space $\mathbb{L}$ of all labels up to time $k$ is given by the disjoint union $\uplus_{\tau\texttt{=}0}^{k}\mathbb{B}_{\tau}$. The state $x$ of an object varies with time, while its label or identity $\ell$ (usually consists of the object’s time of birth and an index to distinguish those born at the same time) is time-invariant. The trajectory of an object is a sequence of consecutive labeled states with a common label.
The cardinality $\left|\bm{X}\right|$ (i.e., number of elements) of the multi-object state $\bm{X}$ varies with time due to the appearance and disappearance of objects. In addition, a multi-object state $\bm{X}$ at any time must have distinct labels. More concisely, let $\mathcal{L}\left(\bm{x}\right)$ denote the label of $\bm{x}$, and for any finite $\bm{X}\subset\mathbb{X}\times\mathbb{L}$, define the label set $\mathcal{L}\left(\bm{X}\right)\triangleq\left\{\mathcal{L}\left(\bm{x}\right):\bm{x}\in\bm{X}\right\}$ and distinct label indicator $\Delta(\bm{X})\triangleq\delta_{|\bm{X}|}\big{[}|\mathcal{L}(\bm{X})|\big{]}$. Then for $\bm{X}$ to be valid multi-object state, we require $\Delta\left(\bm{X}\right)=1$.
In Bayesian estimation, the single-object state and the measurement are modeled as random variables. Hence, for multi-object estimation, the multi-object state and measurement are modeled as Random Finite Sets (RFSs). In the following we describe the so-called standard multi-object models for the dynamics and observations in a multi-object system.
@@ II-A3 Multi-object Observation
Each element $\bm{x}\in\bm{X}$ is either detected with probability $P_{D}(\bm{x})$ and generates an observation $z$ at the sensor with likelihood $g(z|\bm{x})$, or misdetected with probability $1-P_{D}(\bm{x})$. In addition to the detections, the sensor also receives clutter, modeled by a Poisson RFS with intensity function $\kappa$. The multi-object observation $Z$ is the union of detections and clutter. It is assumed that conditional on the multi-object state, objects are detected independently from each other and that clutter and detections are independent.
The association of objects with sensor measurements (at time $k$) is described by a positive 1-1 mapping $\gamma\!:\!\mathbb{L}\!\rightarrow\!\{\texttt{\text{-}}1{\textstyle:}|Z|\}$, i.e., a mapping where no two distinct arguments are mapped to the same positive value [7]. The (extended) association map [^1] Originally called extended association maps, herein referred to as association maps for brevity. $\gamma$ specifies that object (with label) $\ell$ generates measurement $z_{\gamma(\ell)}\in Z$, with $\gamma(\ell)=0$ if it is undetected, and $\gamma(\ell)=-1$ if it does not exist. The positive 1-1 property ensures each measurement comes from at most one object. Let $\mathcal{L}(\gamma)\triangleq\{\ell\in\mathbb{L}:\gamma(\ell)\geq 0\}$ denote the set of live labels [^2] Note the distinction from $\mathcal{L}(\bm{X})$, the labels of a labeled set. of $\gamma$, and $\Gamma$ denote the space of all association maps, then the multi-object likelihood function can be written as [11]
$g(Z|\bm{X})\propto\sum_{\gamma\in\Gamma}\delta_{\mathcal{L}(\gamma)}[\mathcal{L}(\bm{X})][\psi_{Z}^{(\gamma\circ\mathcal{L}(\cdot))}(\cdot)]^{\bm{X}},$ (1)
@@
$\psi_{\{z_{1:M}\}}^{(j)}(\bm{x})=\begin{cases}\frac{P_{D}(\bm{x})g(z_{j}|\bm{x})}{\kappa(z_{j})},&j>0\\ 1-P_{D}(\bm{x}),&j=0\end{cases}.$ (2)